CUHK (SZ) plans to spend ¥50,000 per year to bring in a student from North Korea. The first payment will be made 6 months from today, and the plan will last forever. The discount rate is 7% per year. What is the present value of this commitment?
Answer #1

Present Value of commitment after 6 months =50,000+Payments/Discount Rate =50000+50000/7% =764285.7143
Present Value =764285.7143/(1+7%)^0.5 =738862.89